OpenStudy (anonymous):

At STP, 0.106 g of a gas sample occupies a volume of 62.2 mL. What is the molar mass of this gas? Keep 3 sig figs for your final answer.

OpenStudy (anonymous):

Is it three sig figs or four?

OpenStudy (photon336):

Start with the ideal gas equation; we need number of moles n so we solve for n \[pV = nRT \] then our equation becomes this \[\frac{ pV }{ RT } = n \] \[STP = 1 atm, 298K, 25C = Standard temperature and pressure \] plug everything in \[\frac{ pV }{ RT } = n \frac{ 1atm*0.062L }{ (0.08L, atm, mol, * 298) } = 2.54x10^{-3} mol\] we know that 0.00254 moles contains .106 grams but how many grams are in 1 mole. we can do this by setting up a proportion. \[\frac{ mol }{ x,grams } = \frac{ 0.00254,moles }{ 0.106 grams }\] cross multiply \[0.00254 moles, x = 1.06 grams \] 1 mol contains 44.732 grams \[\frac{ 0.106grams }{ 0.00254mol } = 44.732 grams \]